OpenStudy (anonymous):

Find all polar coordinates of point P where P= -pi/6.

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

P=(1,-pi/6)

OpenStudy (anonymous):

Well, I guess you can just keep adding or subtracting \(2\pi\), so maybe: \[ \forall z\in \mathbb Z\quad \left(1,-\frac \pi6+2\pi z \right) \]

OpenStudy (anonymous):

(1, -pi/6 + (2n + 1)π) or (-1, -pi/6 + 2nπ) (1, -pi/6 + 2nπ) or (-1, -pi/6 + 2nπ) (1, -pi/6 + 2nπ) or (1, -pi/6+ (2n + 1)π) (1, -pi/6 + 2nπ) or (-1, -pi/6 + (2n + 1)π)

OpenStudy (anonymous):

these are my choices

OpenStudy (anonymous):

is it the second one?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

I think it is the last one

OpenStudy (anonymous):

Because it seems: \[ \left(-1,-\frac \pi 6+\pi + 2\pi z\right) \]Are solutions as well, and this simplifies to: \[ \left(-1,-\frac \pi 6+(2 z+1)\pi\right) \]